# Not All Directions Matter: Toward Structured and Task-Aware Low-Rank Adaptation

### 15 Mar 2026 v1

@@ Abstract
Low-Rank Adaptation (LoRA) has become a cornerstone of parameter-efficient fine-tuning (PEFT). Yet, its efficacy is hampered by two fundamental limitations: semantic drift, by treating all update directions with equal importance, and structural incoherence, from adapting layers independently, resulting in suboptimal, uncoordinated updates. To remedy these, we propose StructLoRA, a framework that addresses both limitations through a principled, dual-component design: (1) an Information Bottleneck-guided filter that prunes task-irrelevant directions to mitigate semantic drift, and (2) a lightweight, training-only graph-based coordinator that enforces inter-layer consistency to resolve structural incoherence. Extensive experiments across large language model , vision language model, and vision model (including LLaMA, LLaVA, and ViT) demonstrate that StructLoRA consistently establishes a new state-of-the-art, outperforming not only vanilla LoRA but also advanced dynamic rank allocation and sparsity-based methods. Notably, the benefits are particularly pronounced in challenging low-rank and low-data regimes. Crucially, since our proposed modules operate only during training, StructLoRA enhances performance with zero additional inference cost, advancing the focus of PEFT—from mere parameter compression to a more holistic optimization of information quality and structural integrity.
Not All Directions Matter: Toward Structured and Task-Aware Low-Rank Adaptation
Xi Xiao1†, Chenrui Ma2†, Yunbei Zhang3†, Chen Liu4, Zhuxuanzi Wang1, Yanshu Li5, Lin Zhao6, Guosheng Hu7, Tianyang Wang1, Hao Xu8 1University of Alabama at Birmingham, 2Virginia Tech, 3Tulane University, 4Yale University, 5Brown University, 6Northeastern University, 7University of Bristol, 8Harvard University Correspondence: haxu@bwh.harvard.edu, tw2@uab.edu
@@ 1 Introduction
The advent of large language models (LLMs) has revolutionized natural language processing, yet their immense scale renders full fine-tuning prohibitively expensive for adaptation to downstream tasks. Parameter-Efficient Fine-Tuning (PEFT) has emerged as the de facto solution, enabling customization by updating only a small fraction of a model’s parameters Han et al. (2024). Among these methods, Low-Rank Adaptation (LoRA) Hu et al. (2022a) has become a dominant paradigm due to its elegant simplicity, parameter efficiency, and zero-latency inference Mao et al. (2025). Its success has spurred a family of variants, such as QLoRA Dettmers et al. (2023) and AdaLoRA Zhang et al. (2023b), which further enhance efficiency through quantization and adaptive rank allocation.
@@
However, despite their widespread adoption, we identify two fundamental, unaddressed shortcomings in the LoRA paradigm. The first, semantic drift, stems from allocating a limited parameter budget uniformly across all low-rank update directions, implicitly assuming each direction is equally important for the target task . This assumption overlooks the fact that many directions may capture redundant or noisy signals, leading to inefficient use of capacity and suboptimal performance. The second, structural incoherence, arises from adapting each layer independently, disregarding the inherent compositional structure of deep models like Transformers Raghu et al. (2021); Hu et al. (2022b); Touvron et al. (2022). This layer-wise independence can lead to semantic drift, where update directions across layers become misaligned. We empirically observe this phenomenon, finding low cosine similarity between the update gradients of adjacent layers (e.g., 0.27–0.41), which points to a lack of semantic coordination that can hinder generalization. Sankararaman et al. (2020); Jiang et al. (2024); Carbonnelle and De Vleeschouwer (2018)
@@
To address these critical gaps, we introduce StructLoRA, a novel framework that enhances LoRA with two synergistic components: task-aware filtering and structural coordination. An overview is depicted in Figure 1. To counteract semantic drift, StructLoRA first employs a principled filtering mechanism inspired by the Information Bottleneck principle Tishby and Zaslavsky (2015); Alemi et al. (2017). This module learns to preserve only those update directions that are maximally informative for the task objective by optimizing a constrained objective. Subsequently, to mitigate structural incoherence, we explicitly model inter-layer dependencies by representing the network as a computational graph. A lightweight and tailored Graph Neural Network (GNN) then propagates and refines these filtered updates, encouraging smoother and more aligned adaptation trajectories across the model’s depth. Crucially, this coordination module operates only during training and is discarded at inference, perfectly preserving LoRA’s zero-latency advantage.
@@
We empirically validate the effectiveness of StructLoRA across a comprehensive suite of benchmarks, fine-tuning powerful large language model, vision language model, and vision model such as LLaMA, LLaVA, and ViT on challenging tasks in natural language understanding, computer vision, and multimodal reasoning. Our results demonstrate that StructLoRA consistently and significantly outperforms strong PEFT baselines—including LoRA, AdaLoRA, and DoRA—particularly in low-resource and low-rank regimes. Furthermore, StructLoRA substantially closes the performance gap to full fine-tuning while incurring negligible training overhead and zero additional inference latency.
In summary, our contributions are as follows:
@@
  • To our knowledge, this work is the first to identify and empirically demonstrate the problem of semantic drift and structural incoherence in LoRA-based methods, for which we propose a principled solution that combines task-aware filtering and structural coordination.
@@
  • We propose a new StructLoRA, a lightweight, plug-and-play framework featuring an information bottleneck-guided filter to prune noisy update directions and a graph-based coordinator to enforce inter-layer consistency, which is detached at inference time.
@@
  • We conduct extensive experiments on large language, multimodal and vision models, showing that StructLoRA sets a new state-of-the-art for parameter-efficient fine-tuning, with particularly strong gains in low-rank and few-shot scenarios.
@@ 2.1 The Landscape of PEFT
Adapting large pre-trained models with limited compute has motivated a rich line of Parameter‑Efficient Fine‑Tuning (PEFT) methods Xiao et al. (2026b, 2025b, 2025c); Han et al. (2024); Prottasha et al. (2025); Xin et al. (2024); Braz (2025). Three families are common: additive adapters that insert small trainable blocks Houlsby et al. (2019); He et al. (2022), selective updates that tune only a subset of native weights such as biases Ben-Zaken et al. (2022), and reparameterization approaches that learn a low‑dimensional proxy for the update. Our work follows the third family, whose low‑rank updates merge into the backbone after training and thus add no inference latency. StructLoRA keeps LoRA’s interface but introduces two principles that target overlooked issues in this family: task‑aware directional filtering and cross‑layer structural coordination.
@@ 3.2 Stage 1: Information Bottleneck-Guided Directional Filtering
We learn $\mathbf{m}$ by an information bottleneck (IB) objective Tishby and Zaslavsky (2015); Alemi et al. (2017), as described in Eqn (3), where $\mathcal{L}_{\text{task}}$ is the supervised loss, and $I(\cdot;\cdot)$ denotes mutual information.
@@ 3.4 A Minimal Theoretical View: Coordination as Laplacian Smoothing
In other words, coordination acts like a Laplacian smoother on update directions. Backprop still optimizes $\mathcal{L}_{\text{task}}$; the coordinator adds an explicit structural prior that lowers drift. These quantities are simple to compute and serve as diagnostics together with adjacent-layer cosine in Section 4. For more detailed proof process, please refer to Appendix E.
@@ 4.1 Experimental Setup
Tasks and Datasets
We selected diverse challenging benchmarks to assess performance across different domains. For Natural Language Understanding, we evaluated on the full GLUE benchmark Wang et al. (2018) and a comprehensive set of eight commonsense reasoning benchmarks: BoolQ, PIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, OBQA, and CSQA. Following standard practice, each task was fine-tuned individually on its respective training set. For Natural Language Generation, we used the Magpie-Pro Xu et al. (2024) and OpenPlatypus Lee et al. (2023) datasets to evaluate instruction-following capabilities. For Image Classification, we used ImageNet-1k, CIFAR-100, and the fine-grained Oxford-IIIT Pet dataset. For Multimodal Reasoning, we used established vision-language benchmarks including MS COCO for captioning, VQAv2 and GQA for visual question answering. More details about the baseline, implementation and evaluation can be found in the Appendix A.
Table 1: Main comparison of StructLoRA with baseline PEFT methods. We report primary metrics (Accuracy % for classification/reasoning, CIDEr for captioning) under a comparable budget of trainable parameters ($\sim$0.5–1%). See the Appendix B for additional results.
Method Type BoolQ PIQA CIFAR-100 ImageNet-1k COCO Caption VQAv2
Full Fine-tuning - 82.6 85.3 85.9 78.8 123.5 76.2
Linear Probing - 71.1 74.8 65.7 68.4 105.3 67.0
@@
Method MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg.
@@ 4.2 Main Results
Rank ($r$) Trainable Params (%) BoolQ (LLaMA-7B) CIFAR-100 (ViT-B/16) COCO Caption (LLaVA-1.5-7B)
[…]
Dataset Metric Method 10% Data 25% Data 50% Data 100% Data
BoolQ (LLaMA-7B) Accuracy LoRA 68.5 73.2 76.4 79.1
[…]
CIFAR-100 (ViT-B/16) Accuracy LoRA 73.6 78.0 80.5 81.5
[…]
COCO Caption (LLaVA-1.5-7B) CIDEr LoRA 100.2 108.3 114.0 116.2
[…]
@@
Accuracy (%) $\Delta$ Accuracy (%) $\Delta$ CIDEr $\Delta$
StructLoRA (Full) 81.3 -0.0 84.1 -0.0 122.4 -0.0
@@ 4.3 Ablation Studies
GNN Configuration Accuracy (%)
StructLoRA 81.3
@@
Figure 2: Analysis of filtering strategies. We compare our IB-guided filter with two heuristics under the same keep ratio: Random Masking and Top-$k$ Norm (scored by $\|a_{j}\|_{2}\|b_{j}\|_{2}$ for each rank-one direction). Bars show mean performance across three random seeds; error bars indicate 95% confidence intervals.
@@ E.2 IB-Guided Filtering: Variational Objective
We follow the Information Bottleneck principle: compress nuisance information in $\mathbf{X}$ while preserving label-relevant information about $\mathbf{Y}$ Tishby and Zaslavsky (2015); Alemi et al. (2017). Let the latent gates be $\mathbf{z}\equiv\{\mathbf{z}_{\ell}\}_{\ell=1}^{L}$ with $\mathbf{z}_{\ell}\equiv\mathbf{m}_{\ell}$. A standard variational IB objective is
$\mathcal{L}_{\mathrm{VIB}}\;=\;\mathbb{E}_{q_{\phi}(\mathbf{z}\,|\,\mathbf{X})}\big[-\log p_{\psi}(\mathbf{Y}\,|\,\mathbf{z})\big]\;+\;\beta\;\mathrm{KL}\big(q_{\phi}(\mathbf{z}\,|\,\mathbf{X})\,\|\,r(\mathbf{z})\big),$ (12)
@@ E.3 Depth-Wise Drift Energy and Basic Properties
$\mathcal{E}(\mathbf{U})\;=\;\sum_{\ell=1}^{L-1}\|\mathbf{u}_{\ell+1}-\mathbf{u}_{\ell}\|_{2}^{2}\;=\;\mathbf{U}^{\top}\big(\mathbf{L}\otimes\mathbf{I}\big)\mathbf{U}.$ (13)
@@ E.4 Graph Coordination as Laplacian Smoothing
$\mathbf{U}^{+}\;\approx\;\mathbf{U}+\gamma\,(\mathbf{S}\otimes\mathbf{I})\mathbf{U}\;=\;\big(\mathbf{I}-\gamma\,\mathbf{L}_{\mathrm{norm}}\otimes\mathbf{I}\big)\,\mathbf{U}.$ (14)
@@
$\mathcal{L}_{\mathrm{task}}(\mathbf{U})\;+\;\lambda\,\mathbf{U}^{\top}(\mathbf{L}\otimes\mathbf{I})\mathbf{U}\text{or}$ (16)
@@ E.8 Summary of Verifiable Claims
Task Domain Model Base PEFT Method + StructLoRA Metric Base Performance Enhanced Performance
Language LLaMA-7B QLoRA Dettmers et al. (2023) Accuracy (%) 80.0 81.6 (+1.6)
LLaMA2-7B VeRA Kopiczko et al. (2024) Accuracy (%) 82.6 83.9 (+1.3)
Vision ViT-B/16 LoRA-FA Drost (2024) Accuracy (%) 81.8 83.6 (+1.8)
Multimodal LLaVA-1.5-7B AdapterFusion Pfeiffer et al. (2021) CIDEr 118.3 121.0 (+2.7)

### 25 Apr 2026 v2

# Not All Directions Matter: Towards Structured and Task-Aware Low-Rank Model Adaptation

@@ Abstract
Low-Rank Adaptation (LoRA) has become a cornerstone of parameter-efficient fine-tuning (PEFT). Yet, its efficacy is hampered by two fundamental limitations: semantic drift, arising from treating all update directions with equal importance, and structural incoherence, due to adapting layers independently, resulting in uncoordinated and suboptimal updates. To address these issues, we propose StructLoRA, a framework that tackles both limitations through a principled dual-component design: (1) an Information Bottleneck-guided filter that prunes task-irrelevant directions to mitigate semantic drift, and (2) a lightweight, training-only graph-based coordinator that enforces inter-layer consistency to resolve structural incoherence. Extensive experiments across large language models, vision language models, and vision models (including LLaMA, LLaVA, and ViT) demonstrate that StructLoRA consistently establishes a new state of the art, outperforming not only vanilla LoRA but also advanced dynamic rank allocation and sparsity-based methods. Notably, the gains are particularly pronounced in challenging low-rank and low-data regimes. Crucially, since the proposed modules operate only during training, StructLoRA improves performance with zero additional inference cost, shifting the focus of PEFT from mere parameter compression to a more holistic optimization of information quality and structural integrity.
Not All Directions Matter: Towards Structured and Task-Aware
Low-Rank Model Adaptation
Xi Xiao1thanks: Equal contribution., Chenrui Ma2footnotemark: , Yunbei Zhang3footnotemark: , Chen Liu4, Zhuxuanzi Wang1, Yanshu Li5, Lin Zhao6, Guosheng Hu7, Tianyang Wang1thanks: Corresponding authors., Hao Xu8footnotemark: 1University of Alabama at Birmingham, 2University of Virginia, 3Tulane University, 4Yale University, 5Brown University, 6Northeastern University, 7University of Bristol, 8Harvard University Please direct correspondence to haxu@bwh.harvard.edu or tw2@uab.edu. Project page https://xixiaouab.github.io/StructLoRA/.
@@ 1 Introduction
The advent of large language models (LLMs) has revolutionized natural language processing, yet their immense scale renders full fine-tuning prohibitively expensive for adaptation to downstream tasks. Parameter-Efficient Fine-Tuning (PEFT) has emerged as the de facto solution, enabling customization by updating only a small fraction of a model’s parameters Han et al. (2024). Among these methods, Low-Rank Adaptation (LoRA) Hu et al. (2022a) has become a dominant paradigm due to its simplicity, parameter efficiency, and zero-latency inference Mao et al. (2025). Its success has led to a range of variants, such as QLoRA Dettmers et al. (2023) and AdaLoRA Zhang et al. (2023b), which further improve efficiency through quantization and adaptive rank allocation.
@@
However, despite their widespread adoption, we identify two fundamental and largely unaddressed shortcomings in the LoRA paradigm. The first, semantic drift, stems from allocating a limited parameter budget uniformly across all low-rank update directions, implicitly assuming that each direction is equally important for the target task. This assumption overlooks the fact that many directions may capture redundant or noisy signals, leading to inefficient use of capacity and suboptimal performance. The second, structural incoherence, arises from adapting each layer independently, disregarding the inherent compositional structure of deep models such as Transformers Raghu et al. (2021); Hu et al. (2022b); Touvron et al. (2022); Liu et al. (2026). This layer-wise independence can lead to misaligned update directions across layers. We empirically observe this phenomenon, finding low cosine similarity between the update gradients of adjacent layers (e.g., 0.27–0.41), which indicates a lack of semantic coordination that can hinder generalization Sankararaman et al. (2020); Jiang et al. (2024); Carbonnelle and De Vleeschouwer (2018).
@@
To address these critical gaps, we introduce StructLoRA, a novel framework that enhances LoRA with two synergistic components: task-aware filtering and structural coordination. An overview is shown in Figure 1. To counteract semantic drift, StructLoRA first employs a principled filtering mechanism inspired by the Information Bottleneck principle Tishby and Zaslavsky (2015); Alemi et al. (2017); Liao et al. (2024). This module learns to preserve only those update directions that are maximally informative for the task objective by optimizing a constrained objective.
Subsequently, to mitigate structural incoherence, we explicitly model inter-layer dependencies by representing the network as a computational graph. A lightweight, tailored Graph Neural Network (GNN) then propagates and refines these filtered updates, encouraging smoother and more aligned adaptation trajectories across the model depth. Crucially, this coordination module operates only during training and is discarded at inference, fully preserving LoRA’s zero-latency advantage.
@@
We empirically validate the effectiveness of StructLoRA across a comprehensive suite of benchmarks, fine-tuning powerful large language models, vision language models, and vision models such as LLaMA, LLaVA, and ViT on challenging tasks in natural language understanding, computer vision, and multimodal reasoning. Our results demonstrate that StructLoRA consistently and significantly outperforms strong PEFT baselines, including LoRA, AdaLoRA, and DoRA, particularly in low-resource and low-rank regimes. Furthermore, StructLoRA substantially closes the performance gap to full fine-tuning while incurring negligible training overhead and zero additional inference latency.
In summary, our core contributions include:
@@
  • To our knowledge, this work is the first to identify and empirically demonstrate the problems of semantic drift and structural incoherence in LoRA-based methods, for which we propose a principled solution that combines task-aware filtering and structural coordination.
@@
  • We propose StructLoRA, a lightweight, plug-and-play framework featuring an Information Bottleneck-guided filter to prune noisy update directions and a graph-based coordinator to enforce inter-layer consistency. The coordinator is detached at inference time and therefore incurs no inference overhead.
@@
  • We conduct extensive experiments on large language, multimodal, and vision models, showing that StructLoRA sets a new state of the art for parameter-efficient fine-tuning, with particularly strong gains in low-rank and few-shot scenarios.
@@ 2.1 The Landscape of PEFT
Adapting large pre-trained models with limited compute has motivated a rich line of PEFT methods Xiao et al. (2026b, 2025b, 2025c); Han et al. (2024); Prottasha et al. (2025); Xin et al. (2024); Braz (2025). Three families are common: additive adapters that insert small trainable blocks Houlsby et al. (2019); He et al. (2022), selective updates that tune only a subset of native weights such as biases Ben-Zaken et al. (2022), and reparameterization approaches that learn a low‑dimensional proxy for the update. Our work follows the third family, whose low‑rank updates merge into the backbone after training and thus add no inference latency. StructLoRA keeps LoRA’s interface but introduces two principles that target overlooked issues in this family: task‑aware directional filtering and cross‑layer structural coordination.
@@ 3.2 Stage 1: Information Bottleneck-Guided Directional Filtering
We learn $\mathbf{m}$ by an information bottleneck (IB) objective Tishby and Zaslavsky (2015); Alemi et al. (2017); Liao et al. (2024), as described in Eqn (3), where $\mathcal{L}_{\text{task}}$ is the supervised loss, and $I(\cdot;\cdot)$ denotes mutual information.
@@ 3.4 A Minimal Theoretical View: Coordination as Laplacian Smoothing
In other words, coordination acts like a Laplacian smoother on update directions. Backprop still optimizes $\mathcal{L}_{\text{task}}$; the coordinator adds an explicit structural prior that lowers drift. These quantities are simple to compute and serve as diagnostics together with adjacent-layer cosine in Section 4. For more detailed proofs, please refer to Appendix E.
@@ 4.1 Experimental Setup
Table 1: Main comparison of StructLoRA with baseline PEFT methods. We report primary metrics (Accuracy % for classification, reasoning and question answering; CIDEr for captioning) under a comparable budget of trainable parameters ($\sim$0.5–1%). See Appendix B for additional results.
Method Type Reasoning Image Classification Captioning Visual QA
BoolQ $\uparrow$ PIQA $\uparrow$ CIFAR-100 $\uparrow$ ImageNet-1k $\uparrow$ COCO Caption $\uparrow$ VQAv2 $\uparrow$
Full Fine-tuning 82.6 85.3 85.9 78.8 123.5 76.2
Linear Probing 71.1 74.8 65.7 68.4 105.3 67.0
@@
Method MNLI $\uparrow$ SST-2 $\uparrow$ MRPC $\uparrow$ CoLA $\uparrow$ QNLI $\uparrow$ QQP $\uparrow$ RTE $\uparrow$ STS-B $\uparrow$ Average $\uparrow$
@@
[…]
Tasks and Datasets
We selected diverse challenging benchmarks to assess performance across different domains. For Natural Language Understanding, we evaluated on the full GLUE benchmark Wang et al. (2018) and a comprehensive set of eight commonsense reasoning benchmarks: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-e Clark et al. (2018), ARC-c Clark et al. (2018), OBQA Luo et al. (2022), and CSQA Talmor et al. (2019). Following standard practice, each task was fine-tuned individually on its respective training set. For Natural Language Generation, we used the Magpie-Pro Xu et al. (2024) and OpenPlatypus Lee et al. (2023) datasets to evaluate instruction-following capabilities. For Image Classification, we used ImageNet-1k, CIFAR-100, and the fine-grained Oxford-IIIT Pet dataset. For Multimodal Reasoning, we used established vision-language benchmarks including MS COCO for captioning, VQAv2 and GQA for visual question answering. More details about the baseline, implementation and evaluation can be found in Appendix A.
Rank ($r$) Trainable Params (%) BoolQ $\uparrow$ (LLaMA-7B) CIFAR-100 $\uparrow$ (ViT-B/16) COCO Caption $\uparrow$ (LLaVA-1.5-7B)
[…]
Dataset Metric Method 10% Data 25% Data 50% Data 100% Data
BoolQ (LLaMA-7B) Accuracy $\uparrow$ LoRA 68.5 73.2 76.4 79.1
[…]
CIFAR-100 (ViT-B/16) Accuracy $\uparrow$ LoRA 73.6 78.0 80.5 81.5
[…]
COCO Caption (LLaVA-1.5-7B) CIDEr $\uparrow$ LoRA 100.2 108.3 114.0 116.2
@@ 4.2 Main Results
Accuracy (%) $\uparrow$ $\Delta$ Accuracy (%) $\uparrow$ $\Delta$ CIDEr $\uparrow$ $\Delta$
[…]
@@ 4.3 Ablation Studies
GNN Configuration Accuracy (%) $\uparrow$
[…]
@@
Figure 2: Analysis of filtering strategies. We compare our IB-guided filter with two heuristics under the same keep ratio: Random Masking and Top-$k$ Norm (scored by $\|a_{j}\|_{2}\|b_{j}\|_{2}$ for each rank-one direction). Bars show mean performance across three random seeds.
@@ 5 Conclusion
Impact Statement
Acknowledgments
This manuscript was co-authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.
@@ E.2 IB-Guided Filtering: Variational Objective
We follow the Information Bottleneck principle: compress nuisance information in $\mathbf{X}$ while preserving label-relevant information about $\mathbf{Y}$ Tishby and Zaslavsky (2015); Alemi et al. (2017); Liao et al. (2024). Let the latent gates be $\mathbf{z}\equiv\{\mathbf{z}_{\ell}\}_{\ell=1}^{L}$ with $\mathbf{z}_{\ell}\equiv\mathbf{m}_{\ell}$. A standard variational IB objective is
$\begin{split}\mathcal{L}_{\mathrm{VIB}}=&\;\mathbb{E}_{q_{\phi}(\mathbf{z}\,|\,\mathbf{X})}\big[-\log p_{\psi}(\mathbf{Y}\,|\,\mathbf{z})\big]\\ +&\;\beta\;\mathrm{KL}\big(q_{\phi}(\mathbf{z}\,|\,\mathbf{X})\,\|\,r(\mathbf{z})\big),\end{split}$ (12)
@@ E.3 Depth-Wise Drift Energy and Basic Properties
$\begin{split}\mathcal{E}(\mathbf{U})=&\;\sum_{\ell=1}^{L-1}\|\mathbf{u}_{\ell+1}-\mathbf{u}_{\ell}\|_{2}^{2}\\ =&\;\mathbf{U}^{\top}\big(\mathbf{L}\otimes\mathbf{I}\big)\mathbf{U}.\end{split}$ (13)
@@ E.4 Graph Coordination as Laplacian Smoothing
$\begin{split}\mathbf{U}^{+}\approx&\;\mathbf{U}+\gamma\,(\mathbf{S}\otimes\mathbf{I})\mathbf{U}\\ =&\;\big(\mathbf{I}-\gamma\,\mathbf{L}_{\mathrm{norm}}\otimes\mathbf{I}\big)\,\mathbf{U}.\end{split}$ (14)
@@
$\mathcal{L}_{\mathrm{task}}(\mathbf{U})\;+\;\lambda\,\mathbf{U}^{\top}(\mathbf{L}\otimes\mathbf{I})\mathbf{U}\hskip 4.0pt\text{or}$ (16)
@@ E.8 Summary of Verifiable Claims
Task Domain Model Base PEFT Method Enhancement Method Metric Base Performance Enhanced Performance
Language LLaMA-7B QLoRA Dettmers et al. (2023) StructLoRA Accuracy (%) 80.0 81.6 (+1.6)
LLaMA2-7B VeRA Kopiczko et al. (2024) StructLoRA Accuracy (%) 82.6 83.9 (+1.3)
Vision ViT-B/16 LoRA-FA Drost (2024) StructLoRA Accuracy (%) 81.8 83.6 (+1.8)
Multimodal LLaVA-1.5-7B AdapterFusion Pfeiffer et al. (2021) StructLoRA CIDEr 118.3 121.0 (+2.7)